OpenStudy (anonymous):

Find the sum of the first ten terms of -10 -6 -2 + . . . .

OpenStudy (anonymous):

n sum = ---- [2a + (n-1)d] 2 In your case a = first term = -10 n = number of terms = 10 d = common difference = -6 -(-10) = -6 + 10 = 4 10 so, sum = ---- [2(-10) + (10-1)(4)] = 5[-20 + 36] = 5[16] = 80 2 so sum of first ten terms of the given arithmetic progression is 80